OpenStudy (anonymous):

Help! What is an equation of the line, in point-slope form, that passed through the given point and has the given slope? Point: (8, -8); Slope: 3

OpenStudy (davidusa):

y+8=3x-24

OpenStudy (anonymous):

all of my answer choices end with x-8 or x+8.

OpenStudy (anonymous):

There is a y +8 = 3x-8?

OpenStudy (davidusa):

y+8=3(x-8)